Carbondale Casting produces cast bronze valves on a

12-person

assembly line. On a recent​ day,

180

valves were produced during

a

44​-hour

shift.

​a) Labor productivity of the line​ =

​valves/labor hour ​(round your response to two decimal​ places).

​b) John​ Goodale, the manager at​ Carbondale, changed the layout and was able to increase production to

220

units per

44​-hour

shift. The new labor

productivityequals=

​valves/labor hour ​(round your response to two decimal​ places).

​c) The percentage of productivity

increaseequals=​%

​(round your response to one decimal​ place).
Answer #1

Solution(a)
Given in the question
Total no. of valves produced = 180
No. of hours = 44
Labour productivity of line = 180/44 = 4.09 valves / labour hour
Solution(b)
Total no. of valves produced = 220
Total hours =44
labour productivity = 220/44 = 5 valves/ labour hour
Solution(c)
Increase in productivity percentage = (New productivity - Old Productivity)/Old productivity = (5-4.09)/4.09 = 0.222 or 22.2%